OpenStudy (anonymous):

Suppose y varies directly with x. Write an equation relating x and y. y=-12 when x=20 a y=-2/5x b y=-5/3x c -3/5x d y=3/5x

OpenStudy (lynfran):

y=kx -12=k20 -12/20=k -3/5=k so y=kx would give y=-3/5x

OpenStudy (anonymous):

thank you

OpenStudy (lynfran):

welcome